# A Comprehensive Survey on Continual Learning in Generative Models

### v3

@@ 4.5 Discussions
5 Continual Learning in Vision-Language Action Models
@@
Vision-Language Action (VLA) [218, 219, 20, 220] models refer to exploiting large models to predict action and control behaviors of robots [221, 222] with the aid of a large vision language model. It is of great significance to replace the manipulation of humans in certain risky situations and therefore improve the level of automation. However, robots are inherently adapting to a particular field and inevitably encounter the problem of poor generalization [223] when dealing with the complex real world. Consequently, continual learning in vision language action [124] aims to equip robots with the ability to adapt and acquire new skills in dynamic and ever-changing real-world scenarios like humans do. It is a specific aspect of lifelong robotic learning [224, 225, 226, 227, 228] that employs large multimodal models for robot control and is crucial for robots to accumulate knowledge in novel environments while retaining the ability for known skills to mitigate catastrophic forgetting. Figure 6 illustrates the task formulation of visual-language action models within a continual learning framework.
@@
Due to its wide applicability and potential, many articles have investigated the problem from many aspects based on the general strategy. In the following section, we first introduce the problem setup (Section 5.1) with common training objectives and benchmarks, then we will categorize vision language action methods into three detailed types: replay-based (Section 5.4), architecture-based (Section 5.2) and regularization-based (Section 5.3) approaches, which will be displayed in separate paragraphs.
@@
TABLE III: Summary of continual learning methods for Vision-Language Action Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Tag: I = Image, T = Text, S = Sensor.
@@
Arch Reg Rep
LIBERO [124] LIBERO ViT-T [210], ResNet-T [229], BERT I T Simulated Link
LOTUS [125] LIBERO DINOv2 [230], BERT I T Realistic & Simulated Link
@@
QueST [126] LIBERO, MetaWorld [231] ResNet18, CLIP I T Simulated Link
LEGION [127] LEGION RoBERTa I T Realistic & Simulated Link
CAMA [128] ALFRED [232] ResNet, Bi-LSTM I T Simulated Link
M2Distill [129] LIBERO ResNet, BERT I T S Simulated -
iManip [130] RLBench [233] CLIP, U-Net I T Realistic & Simulated -
RWLA [131] LIBERO R3M [234], SS [235] I T Simulated -
5.1 Problem Setup
Lifelong learning [30, 236] in vision language action extends the concept of the traditional robot action control framework by requiring one robot to continually acquire and adapt across a sequence of tasks $\{T^{1},\dots,T^{K}\}$ over its lifespan while retaining knowledge of previous tasks. This continual learning problem for robots can be formulated as a finite-horizon Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},H,\mu_{0},R)$, where $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function, $H$ is the maximum horizon for each episode of a task, $\mu_{0}$ is the initial state distribution, and $R:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function. During the whole training procedure, each task $T^{k}$ trains the policy $\pi$ from the last task characterized by its own initial state distribution $\mu_{0}^{k}$, while $\mathcal{S},\mathcal{A},\mathcal{T}$ and $H$ remain unchanged across all tasks.
Training objectives. Continual learning in the vision-language action model sequentially trains one single policy $\pi$ through multiple tasks. For each task $T^{k},k\in\{1,\cdots,K\}$, the robot is trained on a dataset of $N$ demonstrations $D^{k}=\{\tau^{k}_{i}\}_{i=1}^{N}$ with a corresponding language description $l^{k}$ guiding the goal of the robot. Each trajectory $\tau^{k}_{i}$ in expert demonstrations with length $L^{k}<H$ consists of state-action pairs $\{(s_{t}^{k},a_{t}^{k})\}_{t=1}^{L^{k}}$ sampled from policy $\pi$ conditioned with task $T^{k}$. The policy is trained with a behavioral cloning loss [237] that aims to mimic the action demonstrations in the dataset:
$\min_{\pi}\mathcal{L}(\pi)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{s_{t}^{k},a_{t}^{k}\sim D^{k}}\left[\sum_{t=1}^{L^{k}}-\log\pi(a_{t}^{k}|s_{\leq t}^{k};T^{k})\right].$ (16)
In the scenario of continual learning, the single policy $\pi$ has to continually adapt to specific requirements of sequential tasks $T^{k}$, while retaining knowledge of gained skills to mitigate catastrophic forgetting.
Benchmarks. To evaluate the performance of robots in continual learning, many benchmarks have been constructed in continual learning scenarios from reinforcement learning to vision language action [238, 239, 124]. For example, Continual World [238] builds a continual reinforcement learning benchmark on robotic manipulation tasks. It provides a realistic, computationally accessible testbed that highlights the limitations of existing continual learning methods in reinforcement learning (RL) settings, which often struggle with forward transfer despite mitigating forgetting. LIBERO [124] builds an imitation learning environment for robots to analyze the continual learning problem of robots from many factors, such as algorithms, architectures, task sequences, variations, and so on, which provides a comprehensive environment for subsequent work.
Figure 6: Illustration of a continual learning framework applied to Vision-Language Action Models.
5.2 Architecture-based Approach
Architecture-based approaches investigate continual learning through structural design, with a focus on enhancing skill representation and reuse. One line of work explores skill-aware architectures that enable models to distinguish, merge, and update skills [240, 126, 125, 127], or construct extensible modular designs for skill composition [241, 242]. LOTUS [125] proposes a hierarchical imitation learning framework that incrementally builds a skill library via unsupervised segmentation of human demonstrations using open-vocabulary vision models. It updates or adds new skills through clustering and composes them via a meta-controller, achieving superior performance in lifelong manipulation through continual discovery and experience replay. IsCiL [243] introduces a prototype-based memory system that incrementally learns shareable skills for adaptation in dynamic environments. By mapping demonstrations to state embeddings and retrieving relevant skill prototypes, it employs adapter modules to preserve prior knowledge. LEGION [127] presents a lifelong reinforcement learning framework that combines Bayesian non-parametric clustering with language-conditioned embeddings. Using a Dirichlet Process Mixture Model (DPMM), it dynamically organizes skills and semantically aligns task understanding, enabling scalable skill accumulation without performance degradation. RLCM [244] formulates a modular lifelong RL framework that composes policies from learned neural modules. It adopts a two-phase learning process—online module selection and adaptation for new tasks, and offline consolidation via batch RL and experience replay—to retain and generalize skills over time.
Another line of research focuses on enhancing robotic action through architectural designs that leverage Large Language Models (LLMs) for high-level guidance in tasks such as planning [245] and skill demonstration [246, 247]. These approaches utilize pre-trained LLMs as knowledge bases, retrieving relevant information via textual prompts to facilitate skill acquisition and enable the transfer of prior knowledge to new tasks. Owing to their data efficiency and rapid inference capabilities, LLMs demonstrate strong potential across diverse robotic scenarios.
5.3 Regularization-based Approach
@@
Regularization methods for continual learning usually try to keep important parameters or features unchanged and impose feature-level or parameter-level regularization during training, therefore maintaining the ability for past skills [248, 128, 129, 243]. CAMA [128] introduces Confidence-Aware Moving Average, a method that dynamically updates stored knowledge (logits) using confidence scores to balance past and current information without requiring task boundaries. PolyTask [248] proposes a method combining task-specific policy training with behavior distillation, which consolidates individual policies into a single model using offline data. With demonstration-guided reinforcement learning and distilling policy outputs, it handles continual actions and lifelong learning by avoiding interference between tasks. M2Distill [129] proposes a multi-modal distillation framework that preserves consistency across vision, language, and action modalities between incremental learning steps. It exploits: (1) multi-modal feature distillation to constrain latent space drift; and (2) policy distillation to maintain alignment of Gaussian mixture model (GMM) action distributions. The approach ensures skill retention without storing prior data, advancing lifelong learning for real-world robotic applications. ELIRL [249] introduces Lifelong Inverse Reinforcement Learning (Lifelong IRL), addressing the challenge of learning multiple sequential tasks and proposing Efficient Lifelong IRL to enable knowledge transfer across tasks, improving performance while minimizing user-provided data.
5.4 Replay-based Approach
@@
Due to the spatial and temporal complexity of robotic manipulation in various and dynamic scenarios, it has become common practice to store part of previous samples and replay them [250, 38] during training procedure to retain the knowledge of previous tasks [130, 127, 241, 131]. iManip [130] addresses the challenge of catastrophic forgetting in robotic manipulation. Given that traditional methods fail to retain previously learned skills, ignoring temporal complexity and action complexity, the introduced framework integrates a temporal replay strategy to preserve task continuity and an extendable PerceiverIO [251] architecture with adaptable action prompts. Experimental results highlight the efficiency, scalability, and robustness across complex, long-horizon tasks. Decision-RWKV [241] exploits a novel sequence modeling framework that integrates the efficiency of the Receptance Weighted Key Value (RWKV) with experience replay mechanisms. It leverages the linear complexity and memory retention capabilities of RWKV to enable efficient lifelong learning across diverse manipulation tasks, and the experience replay combines a replay buffer for historical experiences reuse while acquiring new skills. RWLA [131] combines retrieval-based local adaptation with a selective weighting mechanism to restore forgotten skills. The method retrieves relevant past demonstrations from a memory buffer, identifies critical failure-prone segments, and prioritizes these segments during local fine-tuning, which allows robots to adapt continually while maintaining performance across diverse manipulation tasks.
5.5 Discussions
@@
Continual learning in Vision-Language Action models increasingly adopts hybrid strategies that integrate architectural expansion, replay mechanisms, and regularization constraints. While early methods often apply these components independently, recent advances highlight the importance of balancing them to achieve effective continual learning. Specifically, architectural modifications enable task-specific adaptability, replay mechanisms help preserve cross-task consistency, and regularization mitigates catastrophic interference across modalities. However, key challenges remain, particularly in dynamically coordinating these strategies over time and addressing the tendency of action policy components to forget more rapidly than vision-language alignment modules.
6 Continual Learning in Diffusion Model
@@ 6.1 Problem Setup
Training objectives. Diffusion model is a probabilistic generative model that learns the data distribution by gradually denoising samples drawn from a Gaussian distribution. Specifically, given an initial noise map $\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, a timestep $t\sim\text{Uniform}({1,\dots,T})$, and a user text prompt $p$, the generated image $x$ is obtained as $x=\mathcal{D}(\epsilon_{\theta}(c,t))$, where $c$ is the text embedding extracted from the prompt by a text encoder $\Gamma$ (i.e., $c=\Gamma(p)$). Formally, the diffusion model $\epsilon_{\theta}$ is trained to denoise the noisy image and predict the target as follows:
$\mathcal{L}_{\text{DM}}(\theta)=\mathbb{E}_{z,\varepsilon,c,t}[||\varepsilon-\epsilon_{\theta}(z_{t}|c,t)||^{2}],$ (17)
@@
where $z_{t}$ is the noisy image at timestep $t$, and $\theta$ denotes the parameters of denoising model.
Figure 7: Illustration of a continual learning framework applied to Diffusion Models.
The above process outlines how to generate an image from random noise based on user prompts. To more faithfully capture the user’s intent, PDMs enhance generation by incorporating a small set of user-provided images alongside textual prompts, aiming to bind novel visual concepts (e.g., a specific person or object) to a unique identifier while preserving the model’s generalization ability. To achieve this, DreamBooth [252] introduces a class-specific prior preservation loss, which fine-tunes the model to generate the user-specified subject in various contexts while preventing overfitting and maintaining prior knowledge about the general class. The loss is formulated as:
@@
$\mathcal{L}_{\text{PDM}}(\theta)$ $=\mathbb{E}_{z,\varepsilon,c,t}\left[\left\|\varepsilon-\epsilon_{\theta}(z_{t}|c,t)\right\|^{2}\right]$ (18)
$+\lambda\,\mathbb{E}_{z^{p},\varepsilon,c^{p},t}\left[\left\|\varepsilon-\epsilon_{\theta}(z_{t}^{p}|c^{p},t)\right\|^{2}\right],$
@@
where $z_{t}$ denotes the latent code of an image with the new concept, and $z_{t}^{p}$ corresponds to a class-specific image used for prior preservation. $c$ and $c^{p}$ are the respective conditioning vectors; for instance, $c$ may be “a photo of [$\mathcal{V}^{*}$] dog”, while $c^{p}$ is “a photo of dog”. The special token [$\mathcal{V}^{*}$] is a unique identifier representing the user-specific concept learned from the provided images. $\lambda$ controls the regularization strength.
Continual learning in diffusion models aims to go beyond single-concept personalization by enabling the incremental acquisition of new concepts while avoiding catastrophic forgetting and eliminating the need to retrain on previous data. Specifically, as shown in Figure 7, assume a user intends to continuously generate content based on newly introduced specific concepts. The generation tasks are defined as $\mathcal{T}={\{\mathcal{T}^{k}}\}_{k=1}^{K}$ with corresponding datasets ${\{\mathcal{D}^{k}}\}_{k=1}^{K}$. Each $k$-th generation task consists of $n_{k}$ pairs of concept images $x_{i}^{k}$ and prompts $c_{i}^{k}$, where $K$ denotes the total number of tasks, and $n_{k}$ is typically in the range of 3 to 5 in accordance with existing studies. Accordingly, to learn the $k$-th personalized concept, the model is optimized using the following training loss:
@@
$\mathcal{L}_{\text{CDM}}(\theta)$ $=\mathbb{E}_{z^{k},\varepsilon,c^{k},t}\left[\left\|\varepsilon-\epsilon_{\theta}^{k}(z_{t}^{k}|c^{k},t)\right\|^{2}\right]$ (19)
$+\lambda\,\mathbb{E}_{z^{k,p},\varepsilon,c^{k,p},t}\left[\left\|\varepsilon-\epsilon_{\theta}^{k}(z_{t}^{k,p}|c^{k,p},t)\right\|^{2}\right],$
@@
Backbone. Stable Diffusion (SD-1.5) [22] is a high-resolution image synthesis model based on LDMs, offering a powerful and efficient generative framework. It operates by encoding images into a lower-dimensional latent space using a pre-trained autoencoder, significantly reducing computational costs during training and inference. The model then performs denoising in this latent space with a U-Net backbone guided by a text encoder, such as CLIP, enabling high-fidelity image generation conditioned on text prompts. In continual personalized diffusion settings, LoRA-based fine-tuning is typically applied to the U-Net backbone to efficiently adapt to limited training data.
@@
Benchmark. CLoG [253] proposes two benchmarks: Label-conditioned and Concept-conditional, to evaluate the effectiveness of continual learning approaches in the image generation field. The label-conditioned setting requires the model to learn a sequence of generative tasks conditioned on label indices, while the concept-conditional setting focuses on synthesizing a series of personalized concepts, denoted as $\mathcal{V}^{*}_{k}$, where $k$ indicates the $k$-th concept.
6.2 Architecture-based Approach

### v4

@@ 4.5. Discussions
5. Continual Learning for Vision-Language-Action Models
@@
Vision-Language-Action (VLA) (Chi et al., 2023; Brohan et al., 2023; Kim et al., 2024a; Zawalski et al., 2024) models refer to exploiting large models to predict action and control behaviors of robots (Hansen et al., 2022; López Escoriza et al., 2025) with the aid of a large vision language model. It is of great significance to replace the manipulation of humans in certain risky situations and therefore, improve the level of automation. However, robots are inherently adapting to a particular field and inevitably encounter the problem of poor generalization (Li et al., 2025a) when dealing with the complex real world. Consequently, continual learning in vision language action (Liu et al., 2023b) aims to equip robots with the ability to adapt and acquire new skills in dynamic and ever-changing real-world scenarios as humans do. It is a specific aspect of lifelong robotic learning (Chen et al., 2023a; Gao et al., 2021; Ben-Iwhiwhu et al., 2022; Xie and Finn, 2022; Zhang et al., 2023b) that employs large multimodal models for robot control and is crucial for robots to accumulate knowledge in novel environments while retaining the ability for known skills to mitigate catastrophic forgetting. Figure 6 illustrates the task formulation of visual-language action models within a continual learning framework.
@@
Due to its wide applicability and potential, many articles have investigated the problem from many aspects based on the general strategy. In the following section, we first introduce the problem setup (Section 5.1) with common training objectives and benchmarks, then we will categorize vision language action methods into three detailed types: replay-based (Section 5.4), architecture-based (Section 5.2), and regularization-based (Section 5.3) approaches, which will be displayed in separate paragraphs.
5.1. Problem Setup
Lifelong learning (Kirkpatrick et al., 2017; Yoo et al., 2024) in vision language action extends the concept of the traditional robot action control framework by requiring one robot to continually acquire and adapt across a sequence of tasks $\{T^{1},\dots,T^{K}\}$ over its lifespan while retaining knowledge of previous tasks. This continual learning problem for robots can be formulated as a finite-horizon Markov Decision Process (MDP): $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},H,\mu_{0},R)$, where $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is the transition function, $H$ is the maximum horizon for each episode of a task, $\mu_{0}$ is the initial state distribution, and $R:\mathcal{S}\times\mathcal{A}\rightarrow R$ is the reward function. During the whole training procedure, each task $T^{k}$ trains the policy $\pi$ from the last task characterized by its own initial state distribution $\mu_{0}^{k}$, while $\mathcal{S},\mathcal{A},\mathcal{T}$ and $H$ remain unchanged across all tasks.
Figure 6. Illustration of a continual learning framework applied to Vision-Language-Action Models.
Training objectives. Continual learning in the vision-language-action model sequentially trains one single policy $\pi$ through multiple tasks. For each task $T^{k},k\in\{1,\cdots,K\}$, the robot is trained on a dataset of $N$ demonstrations $D^{k}=\{\tau^{k}_{i}\}_{i=1}^{N}$ with a corresponding language description $l^{k}$ guiding the goal of the robot. Each trajectory $\tau^{k}_{i}$ in expert demonstrations with length $L^{k}<H$ consists of state-action pairs $\{(s_{t}^{k},a_{t}^{k})\}_{t=1}^{L^{k}}$ sampled from policy $\pi$ conditioned with task $T^{k}$. The policy is trained with a behavioral cloning loss (Bain and Sammut, 1995) that aims to mimic the action demonstrations in the dataset:
(15) $\min_{\pi}\mathcal{L}(\pi)=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{s_{t}^{k},a_{t}^{k}\sim D^{k}}\left[\sum_{t=1}^{L^{k}}-\log\pi(a_{t}^{k}|s_{\leq t}^{k};T^{k})\right].$
In the scenario of continual learning, the single policy $\pi$ has to continually adapt to specific requirements of sequential tasks $T^{k}$, while retaining knowledge of gained skills to mitigate forgetting.
Benchmarks. To evaluate the performance of robots in continual learning, many benchmarks have been constructed in continual learning scenarios from reinforcement learning to vision language action (Wołczyk et al., 2021; Powers et al., 2022; Liu et al., 2023b). For example, Continual World (Wołczyk et al., 2021) builds a continual reinforcement learning benchmark on robotic manipulation tasks. It provides a realistic, computationally accessible testbed that highlights the limitations of existing continual learning methods in reinforcement learning settings, which often struggle with forward transfer despite mitigating forgetting. LIBERO (Liu et al., 2023b) builds an imitation learning environment for robots to analyze the continual learning problem of robots from many factors, such as algorithms, architectures, task sequences, and so on, enabling a comprehensive setting for further investigations.
5.2. Architecture-based Approach
Architecture-based continual learning approaches focus on enhancing skill representation and reuse through structural design. One line of work develops skill-aware architectures that distinguish, merge, and update skills (Mete et al., 2024; Wan et al., 2024; Meng et al., 2025) or create modular, extensible frameworks for skill composition (Dong et al., 2025; Rauso et al., 2024). For example, LOTUS (Wan et al., 2024) incrementally builds a skill library by unsupervisedly segmenting human demonstrations with open-vocabulary vision models, updating skills via clustering, and composing them with a meta-controller for lifelong manipulation. IsCiL (Lee et al., 2024) employs a prototype-based memory to learn shareable skills, retrieving relevant prototypes via state embeddings and preserving knowledge with adapters. LEGION (Meng et al., 2025) combines Bayesian non-parametric clustering and language-conditioned embeddings, dynamically organizing skills with a Dirichlet Process Mixture Model to scale skill accumulation without performance loss. RLCM (Mendez et al., 2022) presents a modular lifelong RL framework that dynamically composes policies from reusable neural modules. It combines online module selection and adaptation with offline consolidation via batch RL and experience replay to ensure long-term skill retention and generalization. Jia et al. (Jia et al., 2025) propose hierarchical routing with cross-modal clustering and SVD-based incremental LoRA, preserving principal components and orthogonally training residuals to reduce catastrophic forgetting in embodied continual learning. DRAE (Long et al., 2025) integrates dynamic MoE routing with parameterized RAG and hierarchical RL to enable lifelong learning while mitigating catastrophic forgetting through spectral-aware consolidation and DPMM-based knowledge retention.
Another line leverages Large Language Models (LLMs) for high-level task guidance in robotics, such as planning (Parakh et al., 2024) and skill demonstration (Tziafas and Kasaei, 2024; Bärmann et al., 2024). These methods utilize pre-trained LLMs as knowledge bases, retrieving task-relevant information via textual prompts to aid skill acquisition and transfer. Thanks to their data efficiency and fast inference, LLMs show great promise in diverse robotic continual learning scenarios.
@@ 5.2. Architecture-based Approach
Table 3. Summary of continual learning methods for Vision-Language-Action Models. Taxonomy: indicates the primary solution, and ✔ indicates the supporting method. Tag: I = Image, T = Text, S = Sensor.
@@
Arch. Reg. Rep.
LIBERO (Liu et al., 2023b) LIBERO ViT-T, ResNet-T, BERT I T Simulated Link
LOTUS (Wan et al., 2024) LIBERO DINOv2, BERT I T Realistic & Simulated Link
@@
QueST (Mete et al., 2024) LIBERO, MetaWorld ResNet18, CLIP I T Simulated Link
LEGION (Meng et al., 2025) LEGION RoBERTa I T Realistic & Simulated Link
Jia et al. (Jia et al., 2025) ALFRED LLAMA-2-7B, CLIP I T Simulated -
DRAE (Long et al., 2025) MimicGen, NAVSIM ResNet I T Simulated -
CAMA (Kim et al., 2024b) ALFRED ResNet, Bi-LSTM I T Simulated Link
M2Distill (Roy et al., 2024) LIBERO ResNet, BERT I T S Simulated -
iManip (Zheng et al., 2025c) RLBench CLIP, U-Net I T Realistic & Simulated -
RWLA (Yang et al., 2024b) LIBERO R3M, SS I T Simulated -
5.3. Regularization-based Approach
@@
Regularization methods for continual learning usually try to keep important parameters or features unchanged and impose feature-level or parameter-level regularization during training, therefore maintaining the ability for past skills (Haldar and Pinto, 2023; Kim et al., 2024b; Roy et al., 2024; Lee et al., 2024). CAMA (Kim et al., 2024b) introduces Confidence-Aware Moving Average, a method that dynamically updates stored knowledge (logits) using confidence scores to balance past and current information without requiring task boundaries. PolyTask (Haldar and Pinto, 2023) proposes a method combining task-specific policy training with behavior distillation, which consolidates individual policies into a single model using offline data. With demonstration-guided reinforcement learning and distilling policy outputs, it handles continual actions and lifelong learning by avoiding interference between tasks. M2Distill (Roy et al., 2024) proposes a multi-modal distillation framework that preserves consistency across vision, language, and action modalities between incremental learning steps. It exploits: (1) multi-modal feature distillation to constrain latent space drift; and (2) policy distillation to maintain alignment of Gaussian mixture model (GMM) action distributions. The approach ensures skill retention without storing prior data, advancing lifelong learning for real-world robotic applications. ELIRL (Mendez et al., 2018) introduces Lifelong Inverse Reinforcement Learning (Lifelong IRL), addressing the challenge of learning multiple sequential tasks and proposing Efficient Lifelong IRL to enable knowledge transfer across tasks, improving performance while minimizing user-provided data.
5.4. Replay-based Approach
@@
Due to the spatial and temporal complexity of robotic manipulation in various and dynamic scenarios, it has become common practice to store part of previous samples and replay them (Bang et al., 2021; Rolnick et al., 2019) during the training procedure to retain the knowledge of previous tasks (Zheng et al., 2025c; Meng et al., 2025; Dong et al., 2025; Yang et al., 2024b). iManip (Zheng et al., 2025c) addresses the challenge of catastrophic forgetting in robotic manipulation. Given that traditional methods fail to retain previously learned skills, ignoring temporal complexity and action complexity, the introduced framework integrates a temporal replay strategy to preserve task continuity and an extendable PerceiverIO (Jaegle et al., 2021) architecture with adaptable action prompts. Experimental results highlight the efficiency, scalability, and robustness across complex, long-horizon tasks. Decision-RWKV (Dong et al., 2025) exploits a novel sequence modeling framework that integrates the efficiency of the Receptance Weighted Key Value (RWKV) with experience replay mechanisms. It leverages the linear complexity and memory retention capabilities of RWKV to enable efficient lifelong learning across diverse manipulation tasks, and the experience replay combines a replay buffer for historical experiences reuse while acquiring new skills. RWLA (Yang et al., 2024b) combines retrieval-based local adaptation with a selective weighting mechanism to restore forgotten skills. The method retrieves relevant past demonstrations from a memory buffer, identifies critical failure-prone segments, and prioritizes these segments during local fine-tuning, which allows robots to adapt continually while maintaining performance across diverse manipulation tasks.
5.5. Discussions
@@
Continual learning in Vision-Language-Action models increasingly adopts hybrid strategies that integrate architectural expansion, replay mechanisms, and regularization constraints. While early methods often apply these components independently, recent advances highlight the importance of balancing them to achieve effective continual learning. Specifically, architectural modifications enable task-specific adaptability, replay mechanisms help preserve cross-task consistency, and regularization mitigates catastrophic interference across modalities. However, key challenges remain, particularly in dynamically coordinating these strategies over time and addressing the tendency of action policy components to forget more rapidly than vision-language alignment modules.
6. Continual Learning for Diffusion Models
@@ 6.1. Problem Setup
Training objectives. Diffusion models are probabilistic generative models that learn data distributions by iteratively denoising Gaussian noise. Given an initial noise map $\varepsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, timestep $t\sim\text{Uniform}({1,\dots,T})$, and user prompt $p$, an image is generated as $x=\mathcal{D}(\epsilon_{\theta}(c,t))$, where $c=\Gamma(p)$ is the text embedding from encoder $\Gamma$. The denoising model $\epsilon_{\theta}$ is trained with:
[…]
@@
where $z_{t}$ is the noisy image at timestep $t$, and $\theta$ denotes the parameters of the denoising model.
To better reflect user intent, PDMs enhance generation by incorporating a few user images alongside text prompts to bind novel concepts (e.g., a person or object) to a unique identifier while retaining generalization. DreamBooth (Ruiz et al., 2023) introduces a class-specific prior preservation loss to prevent overfitting and preserve base class knowledge:
@@
(17) $\mathcal{L}_{\text{PDM}}(\theta)=\mathbb{E}_{z,\varepsilon,c,t}\left[\left\|\varepsilon-\epsilon_{\theta}(z_{t},c,t)\right\|^{2}\right]+\lambda\,\mathbb{E}_{z^{p},\varepsilon,c^{p},t}\left[\left\|\varepsilon-\epsilon_{\theta}(z_{t}^{p},c^{p},t)\right\|^{2}\right],$
@@
where $z_{t}$ denotes the latent code of an image with the new concept, and $z_{t}^{p}$ corresponds to a class-specific image used for prior preservation. $c$ and $c^{p}$ are the respective conditioning vectors; for instance, $c$ may be “a photo of $\mathcal{V}^{*}$] dog”, while $c^{p}$ is “a photo of dog”. The special token $\mathcal{V}^{*}$] is a unique identifier representing the user-specific concept learned from the provided images. $\lambda$ controls the regularization strength.
Figure 7. Illustration of a continual learning framework applied to Diffusion Models.
Continual learning in diffusion models extends beyond single-concept personalization by incrementally acquiring new concepts without retraining on prior data or suffering catastrophic forgetting. As illustrated in Figure 7, let $\mathcal{T}={\{\mathcal{T}^{k}}\}_{k=1}^{K}$ be a sequence of $K$ generation tasks with datasets ${\{\mathcal{D}^{k}}\}_{k=1}^{K}$. Each task $\mathcal{T}^{k}$ contains $n_{k}$ image-prompt pairs $(x_{i}^{k},c_{i}^{k})$, where $n_{k}$ is typically 3-5. The model learns the $k$-th concept via the following loss:
@@
(18) $\mathcal{L}_{\text{CDM}}(\theta)=\mathbb{E}_{z^{k},\varepsilon,c^{k},t}\left[\left\|\varepsilon-\epsilon_{\theta}^{k}(z_{t}^{k},c^{k},t)\right\|^{2}\right]+\lambda\,\mathbb{E}_{z^{k,p},\varepsilon,c^{k,p},t}\left[\left\|\varepsilon-\epsilon_{\theta}^{k}(z_{t}^{k,p},c^{k,p},t)\right\|^{2}\right],$
@@
Backbone. Stable Diffusion (SD-1.5) (Rombach et al., 2022) is a high-resolution image synthesis model based on LDMs, offering a powerful and efficient generative framework. It operates by encoding images into a lower-dimensional latent space using a pre-trained autoencoder, significantly reducing computational costs during training and inference. The model then performs denoising in this latent space with a U-Net backbone guided by a text encoder, such as CLIP, enabling high-fidelity image generation conditioned on text prompts. In continual personalized diffusion settings, LoRA-based fine-tuning is typically applied to the U-Net backbone to efficiently adapt to limited training data.
@@
Benchmark. CLoG (Zhang et al., 2024d) proposes two benchmarks: Label-conditioned and Concept-conditional, to evaluate the effectiveness of continual learning approaches in the image generation field. The label-conditioned setting requires the model to learn a sequence of generative tasks conditioned on label indices, while the concept-conditional setting focuses on synthesizing a series of personalized concepts, denoted as $\mathcal{V}^{*}_{k}$, where $k$ indicates the $k$-th concept.
6.2. Architecture-based Approach